# Deep Learning for Code in the Agentic Era

### v1

@@ 1 Introduction
Modern code-capable language models increasingly interact with repositories by both analyzing and generating code diffs. In tasks such as issue resolution [5; 17], CI build repair [6], or bug repair [13], the model’s task is to produce a patch for an existing codebase that is then evaluated for correctness. Tasks like commit message generation treat diffs as an input that a model should analyze to generate a summary [11; 29].
There exists a multitude of formats for representing code diffs. Classical diff formats include the (i) normal format – changes without surrounding context, (ii) context format – changes with surrounding lines, and (iii) unified format [^1] Throughout the work, we use unified diff or udiff for all the diffs that use the unified diff syntax. – a compact, context-carrying variant with @@ hunk headers and +/- prefixes (see the example in Figure 6). These three formats are standardized in GNU diffutils [21] and remain the basis for machine-applicable patches. In modern tooling, git diff produces a Git patch similar to unified diff but with additional Git-specific headers.
While benchmarks such as SWE-bench evaluate patches in the unified diff format, the choice of the diff representation used by an LLM can influence the quality of generations and their cost. For example, Aider reports that switching from a search/replace scheme to unified diffs cuts unhelpful outputs and raises benchmark scores [1]. OpenAI trains GPT-4.1 to generate and apply diffs in a recommended V4A patch format, and reports large gains on Aider’s polyglot diff benchmark [2; 23]. Existing works [25; 27; 28] report using different approaches, such as working with unified diffs, performing rewrites on specified ranges, supporting search-replace formats that specify anchored replacements.
Despite this centrality, current evaluations make it hard to isolate the effect of diff format from other factors like retrieval or tool use. Wang et al. [26] analyze SWE-bench and highlight that there exist distinct failure modes: a patch may fail because it is syntactically malformed, because it does not apply to the intended context, or because it applies but does not fix the issue. This shows that patch outcomes depend on more than just representation, making it difficult to isolate and evaluate the impact of diff formats within such end-to-end benchmarks.
@@
First, we apply Diff-XYZ to analyze how well proprietary and open-source LLMs work with the most widespread unified diff format. Proprietary models consistently outperform open-source models, with Claude 4 Sonnet and GPT-4.1 achieving the highest scores across the tasks; GPT-4.1 is more sensitive to a task prompt and tends to emit V4A diff format by default. For open models, scores improve as model size grows, yet performance is still short of strong results, especially for smaller open models. Diff-XYZ can help by providing focused, reproducible checks that highlight where formats and prompts make the most difference. This analysis clarifies model behavior on the benchmark tasks and establishes strong reference results. Building on these findings, we then compare alternative edit representations: udiff-based formats work best for Apply and Anti-Apply, search-replace excels for Diff Generation, but for smaller models modified udiff variants perform best.
@@ 3.1 Apply Task and Anti-Apply Task
Apply is a straightforward task: given the old code and the diff, generate the new code. It can be viewed as a special case of code generation task with all the needed information being already in the prompt, yet the model has to correctly interpret the diff to solve it. If an LLM is capable of understanding the diff and copying, we expect it to solve the task accurately, with larger models solving it near-perfectly. In case a model scores low on the Apply task, it may require additional fine-tuning to work with diffs or an adaptation of the diff format. Note that Apply is a natural simplification of the recently introduced problem of smart paste or fuzzy diff application with LLMs [12], i.e., generate a new code state after applying a patch in some insufficient format.
@@ 3.2 Diff Generation Task
Unlike Apply and Anti-Apply, diff calculation is not uniquely defined. Even for the same code change, different but valid diffs can be produced, depending on factors such as the number of context lines. Because of that, we cannot apply similarity metrics like EM or IoU to the Diff Generation task. To overcome this problem, we evaluate EM and IoU for the new code and the code that is the result of applying the generated diff to the old code. Because diff application is a strict, all-or-nothing procedure, EM and IoU after application only capture exact correctness when the generated diff applies successfully. To measure partial correctness even when application fails, we also compute F1-scores over the sets of added and deleted lines [20].
@@ 4.1 Proprietary Models
Proprietary LLMs are usually more powerful than open-source ones [3; 10; 16; 23], and since the tasks are conceptually simple, we expect the best models to achieve near-perfect scores.
@@
GPT 4o w/o format 0.34 0.50 0.72 0.51 0.52 0.89
GPT 4o w/format 0.33 0.53 0.79 0.52 0.55 0.98
GPT 4o-mini w/o format 0.17 0.38 0.65 0.36 0.40 0.92
GPT 4o-mini w/format 0.16 0.27 0.60 0.33 0.28 0.90
GPT 4.1 w/o format 0.34 0.35 0.34 0.25 0.35 0.44
GPT 4.1 w/format 0.76 0.78 0.76 0.52 0.79 0.99
GPT 4.1 mini w/o format 0.11 0.12 0.08 0.08 0.12 0.15
GPT 4.1 mini w/format 0.60 0.64 0.52 0.33 0.78 0.65
GPT 4.1 nano w/o format 0.25 0.34 0.26 0.14 0.35 0.47
GPT 4.1 nano w/format 0.21 0.27 0.18 0.09 0.28 0.53
Claude 4 Sonnet w/o format 0.64 0.66 0.73 0.57 0.67 0.80
Claude 4 Sonnet w/format 0.85 0.89 0.92 0.71 0.89 1.00
Gemini 2.5 Flash w/o format 0.69 0.79 0.82 0.61 0.80 0.93
Gemini 2.5 Flash w/format 0.66 0.76 0.80 0.59 0.77 0.90
@@ 4.2 Open Source Models
To analyze how model size affects diff understanding, we evaluate the Qwen2.5-Coder series [15], a family of open-source, code-focused LLMs ranging from 0.5B to 32B parameters. Qwen2.5-Coder models achieve near state-of-the-art results on standard code generation and reasoning benchmarks among open-source models, making them a strong testbed for scaling analysis on Diff-XYZ. Tables 3 and 4 report results for the Apply/Anti-Apply and Diff Generation tasks, respectively.
@@
Qwen2.5-Coder-0.5B-Instruct 0.00 0.00 0.01 0.02 0.01 0.23
Qwen2.5-Coder-1.5B-Instruct 0.01 0.04 0.07 0.10 0.04 0.69
Qwen2.5-Coder-3B-Instruct 0.00 0.12 0.18 0.15 0.15 0.72
Qwen2.5-Coder-7B-Instruct 0.03 0.17 0.29 0.24 0.19 0.70
Qwen2.5-Coder-14B-Instruct 0.14 0.35 0.47 0.37 0.38 0.92
Qwen2.5-Coder-32B-Instruct 0.24 0.47 0.61 0.50 0.50 0.99
@@ 5 Diff Format Exploration
    search-replace: a sequence of edits where each search substring is replaced by a replace substring, following Aider [1] and Li et al. [18] (Figure 9).
@@
GPT 4o udiff 0.86 0.85 0.43 0.89 0.83
udiff-h 0.85 0.89 0.03 0.20 0.61
udiff-l 0.82 0.88 0.03 0.11 0.25
search-replace 0.57 0.60 0.74 0.93 0.89
GPT 4.1 udiff 0.90 0.88 0.77 0.92 0.91
udiff-h 0.92 0.93 0.06 0.25 0.61
udiff-l 0.91 0.88 0.06 0.21 0.39
search-replace 0.57 0.56 0.95 0.97 0.94
GPT 4.1-mini udiff 0.89 0.81 0.73 0.87 0.84
udiff-h 0.86 0.85 0.03 0.28 0.62
udiff-l 0.86 0.83 0.04 0.10 0.14
search-replace 0.57 0.55 0.90 0.94 0.94
GPT 4.1-nano udiff 0.37 0.02 0.51 0.81 0.78
udiff-h 0.32 0.01 0.12 0.47 0.49
udiff-l 0.44 0.04 0.09 0.10 0.11
search-replace 0.33 0.01 0.05 0.07 0.07
Claude 4 Sonnet udiff 0.95 0.82 0.82 0.95 0.94
udiff-h 0.95 0.87 0.02 0.13 0.31
udiff-l 0.95 0.93 0.15 0.34 0.55
search-replace 0.57 0.48 0.94 0.97 0.96
Gemini 2.5 Flash udiff 0.90 0.34 0.73 0.94 0.92
udiff-h 0.91 0.54 0.10 0.22 0.28
udiff-l 0.79 0.07 0.17 0.32 0.26
search-replace 0.57 0.53 0.88 0.96 0.95
@@
udiff-h 0.01 0.01 0.03 0.18 0.18
udiff-l 0.00 0.01 0.23 0.57 0.42
search-replace 0.00 0.00 0.00 0.04 0.22
Qwen2.5-Coder-1.5B udiff 0.16 0.04 0.01 0.07 0.09
udiff-h 0.16 0.05 0.20 0.45 0.42
udiff-l 0.11 0.03 0.38 0.49 0.49
search-replace 0.20 0.07 0.19 0.35 0.38
Qwen2.5-Coder-3B udiff 0.38 0.17 0.01 0.20 0.22
udiff-h 0.41 0.20 0.02 0.16 0.21
udiff-l 0.34 0.13 0.36 0.54 0.55
search-replace 0.28 0.25 0.14 0.30 0.22
Qwen2.5-Coder-7B udiff 0.58 0.65 0.05 0.39 0.31
udiff-h 0.57 0.64 0.01 0.17 0.28
udiff-l 0.57 0.62 0.07 0.17 0.18
search-replace 0.50 0.50 0.28 0.62 0.58
Qwen2.5-Coder-32B udiff 0.84 0.87 0.23 0.73 0.63
udiff-h 0.83 0.86 0.00 0.25 0.67
udiff-l 0.83 0.82 0.03 0.18 0.39
search-replace 0.57 0.53 0.68 0.92 0.86
The results are presented in Table 5. For Diff Generation, search-replace is a strong default for most larger models, while udiff-l achieves the best scores for smaller models, a result that is both consistent and unexpected. Another unexpected finding is that udiff-h is typically much worse than udiff, despite only relaxing the hunk header and making a minimal change in format. For Apply and Anti-Apply, however, search-replace underperforms relative to more structured formats, highlighting a trade-off between ease of generating edits and faithfulness of application. These results suggest that while search–replace is attractive for generation, it is a poor fit for tasks where faithful application matters. This may also affect downstream tasks such as commit message generation, which rely on diffs as structured inputs. For such tasks, explicit udiff variants remain the safer and more reliable choice.
@@
    Marker collisions. Small models may confuse single-character udiff markers (+, -, leading space) with ordinary code characters. Replacing them with explicit tags (ADD/DEL/CON) in udiff-l makes the control tokens unambiguous and rare; the model mainly needs to decide, for each line, which tag to use and then emit the line, an easier decision for weaker models.
@@ 7 Related Work
A long line of benchmarks evaluate LLMs on code generation from natural language. Chen et al. [9] and Austin et al. [4] introduced HumanEval and MBPP, which remain the most widely used, with extensions such as HumanEval+ and MBPP+ [19], HumanEval-XL [24], and multilingual variations such as MultiPL-E [7]. BigCodeBench [30] emphasizes more complex and library-heavy tasks. These benchmarks primarily test function-level synthesis and correctness, but do not address how edits are represented or applied.
@@
More recent work has evaluated language models in editing settings. Cassano et al. [8] introduces instruction-following edits, and Guo et al. [14] evaluate debugging, polishing and translation tasks. At a larger scale, Jimenez et al. [17] proposed SWE-bench, which tasks models with resolving real GitHub issues by producing patches that apply and pass tests. These benchmarks reflect realistic workflows, but include many factors, from retrieval and long context reasoning to semantic correctness and patch formatting. These benchmarks capture realistic workflows, including patch application, but their evaluation mixes retrieval, long-context reasoning, semantic correctness, and patch formatting — making it difficult to isolate the role of edit representation.
@@
Despite this breadth, existing benchmarks all assume a fixed diff format, typically unified diff. In reality, there are multiple representations in use: normal diff, context diff, unified diff, OpenAI’s V4A diff format [23], and search/replace schemes, including anchored replacements used by some agents. Format is therefore a real variable, not a constant. Aider’s benchmarks [2] are notable for reporting differences between formats (whole file rewrite, search/replace, unified diff), but their evaluation was motivated by tool design choices rather than a systematic study of edit representations.

### v2

@@ 1 Introduction
Modern code-capable language models increasingly interact with repositories by both analyzing and generating code diffs. In tasks such as issue resolution [swe-rebench; swebench], CI build repair [LCA], or bug repair [defects4j], the model’s task is to produce a patch for an existing codebase that is then evaluated for correctness. Tasks like commit message generation treat diffs as an input that a model should analyze to generate a summary [CMG; CMG-2].
There exists a multitude of formats for representing code diffs. Classical diff formats include the (i) normal format – changes without surrounding context, (ii) context format – changes with surrounding lines, and (iii) unified format [^1] Throughout the work, we use unified diff or udiff for all the diffs that use the unified diff syntax. – a compact, context-carrying variant with @@ hunk headers and +/- prefixes (see the example in Figure 6). These three formats are standardized in GNU diffutils [diffutils] and remain the basis for machine-applicable patches. In modern tooling, git diff produces a Git patch similar to unified diff but with additional Git-specific headers.
While benchmarks such as SWE-bench evaluate patches in the unified diff format, the choice of the diff representation used by an LLM can influence the quality of generations and their cost. For example, Aider reports that switching from a search/replace scheme to unified diffs cuts unhelpful outputs and raises benchmark scores [Aider-udiff]. OpenAI trains GPT-4.1 to generate and apply diffs in a recommended V4A patch format, and reports large gains on Aider’s polyglot diff benchmark [aider-polyglot; gpt4.1]. Existing works [refactai; demystifying; swe-agent] report using different approaches, such as working with unified diffs, performing rewrites on specified ranges, supporting search-replace formats that specify anchored replacements.
Despite this centrality, current evaluations make it hard to isolate the effect of diff format from other factors like retrieval or tool use. wang2025solved analyze SWE-bench and highlight that there exist distinct failure modes: a patch may fail because it is syntactically malformed, because it does not apply to the intended context, or because it applies but does not fix the issue. This shows that patch outcomes depend on more than just representation, making it difficult to isolate and evaluate the impact of diff formats within such end-to-end benchmarks.
@@
First, we apply Diff-XYZ to analyze how well proprietary and open-source LLMs work with the most widespread unified diff format. Proprietary models consistently outperform open-source models, with Claude 4 Sonnet and GPT-4.1 achieving the highest scores across the tasks; GPT-4.1 is more sensitive to a task prompt and tends to emit V4A diff format by default. For open models, scores improve as model size grows, yet performance is still short of strong results, especially for smaller open models. Diff-XYZ can help by providing focused, reproducible checks that highlight where formats and prompts make the most difference. This analysis clarifies model behavior on the benchmark tasks and establishes strong reference results. Building on these findings, we then compare alternative edit representations: search-replace format consistently performs best across all tasks for stronger models, while udiff and udiff-h remain competitive alternatives. For smaller open models, performance remains limited regardless of format.
@@ 3.1 Apply Task and Anti-Apply Task
Apply is a straightforward task: given the old code and the diff, generate the new code. It can be viewed as a special case of code generation task with all the needed information being already in the prompt, yet the model has to correctly interpret the diff to solve it. If an LLM is capable of understanding the diff and copying, we expect it to solve the task accurately, with larger models solving it near-perfectly. In case a model scores low on the Apply task, it may require additional fine-tuning to work with diffs or an adaptation of the diff format. Note that Apply is a natural simplification of the recently introduced problem of smart paste or fuzzy diff application with LLMs [smart-paste], i.e., generate a new code state after applying a patch in some insufficient format.
@@ 3.2 Diff Generation Task
Unlike Apply and Anti-Apply, diff calculation is not uniquely defined. Even for the same code change, different but valid diffs can be produced, depending on factors such as the number of context lines. Because of that, we cannot apply similarity metrics like EM or IoU to the Diff Generation task. To overcome this problem, we evaluate EM and IoU for the new code and the code that is the result of applying the generated diff to the old code. Because diff application is a strict, all-or-nothing procedure, EM and IoU after application only capture exact correctness when the generated diff applies successfully. To measure partial correctness even when application fails, we also compute F1-scores over the sets of added and deleted lines [info-retrieval].
@@ 4.1 Proprietary Models
Proprietary LLMs are usually more powerful than open-source ones [claude; gemini; gpt-4o; gpt4.1], and since the tasks are conceptually simple, we expect the best models to achieve near-perfect scores.
@@
GPT 4o w/o format 0.37 0.52 0.74 0.71 0.53 0.83
GPT 4o w/format 0.38 0.56 0.86 0.80 0.58 0.97
GPT 4o-mini w/o format 0.18 0.38 0.70 0.60 0.40 0.93
GPT 4o-mini w/format 0.17 0.28 0.66 0.56 0.29 0.91
GPT 4.1 w/o format 0.34 0.36 0.42 0.41 0.36 0.43
GPT 4.1 w/format 0.76 0.78 0.96 0.96 0.79 0.99
GPT 4.1-mini w/o format 0.12 0.12 0.15 0.14 0.13 0.16
GPT 4.1-mini w/format 0.60 0.64 0.74 0.72 0.64 0.77
GPT 4.1-nano w/o format 0.23 0.31 0.38 0.37 0.32 0.45
GPT 4.1-nano w/format 0.21 0.27 0.41 0.38 0.28 0.54
Claude 4 Sonnet w/o format 0.62 0.65 0.76 0.75 0.65 0.79
Claude 4 Sonnet w/format 0.85 0.89 0.96 0.95 0.89 1.00
Gemini 2.5 Flash w/o format 0.75 0.86 0.93 0.92 0.87 0.99
Gemini 2.5 Flash w/format 0.71 0.82 0.95 0.92 0.84 1.00
@@ 4.2 Open Source Models
To analyze how model size affects diff understanding, we evaluate the Qwen2.5-Coder series [qwencoder], a family of open-source, code-focused LLMs ranging from 0.5B to 32B parameters. Qwen2.5-Coder models achieve near state-of-the-art results on standard code generation and reasoning benchmarks among open-source models, making them a strong testbed for scaling analysis on Diff-XYZ. Tables 3 and 4 report results for the Apply/Anti-Apply and Diff Generation tasks, respectively.
@@
Qwen2.5-Coder-0.5B-Instruct 0.00 0.00 0.00 0.02 0.01 0.23
Qwen2.5-Coder-1.5B-Instruct 0.01 0.04 0.16 0.18 0.04 0.70
Qwen2.5-Coder-3B-Instruct 0.00 0.04 0.26 0.22 0.06 0.84
Qwen2.5-Coder-7B-Instruct 0.03 0.17 0.32 0.27 0.19 0.70
Qwen2.5-Coder-14B-Instruct 0.14 0.35 0.63 0.53 0.38 0.92
Qwen2.5-Coder-32B-Instruct 0.23 0.46 0.77 0.65 0.49 0.99
@@ 5 Diff Format Exploration
    search-replace: a sequence of edits where each search substring is replaced by a replace substring, following Aider-udiff and patchpilot (Figure 9).
@@
GPT 4o udiff 0.86 0.85 0.41 0.90 0.84
udiff-h 0.85 0.89 0.30 0.80 0.72
udiff-l 0.82 0.88 0.01 0.02 0.02
search-replace 0.94 0.94 0.73 0.91 0.83
GPT 4.1 udiff 0.90 0.88 0.81 0.97 0.97
udiff-h 0.92 0.93 0.69 0.80 0.79
udiff-l 0.91 0.88 0.08 0.11 0.10
search-replace 0.96 0.93 0.95 0.97 0.93
GPT 4.1-mini udiff 0.89 0.81 0.78 0.90 0.88
udiff-h 0.86 0.85 0.69 0.84 0.85
udiff-l 0.86 0.83 0.01 0.02 0.02
search-replace 0.94 0.84 0.92 0.95 0.93
GPT 4.1-nano udiff 0.37 0.02 0.50 0.82 0.78
udiff-h 0.32 0.01 0.44 0.79 0.75
udiff-l 0.44 0.04 0.00 0.00 0.00
search-replace 0.54 0.01 0.07 0.08 0.08
Claude 4 Sonnet udiff 0.95 0.82 0.82 0.97 0.95
udiff-h 0.95 0.87 0.83 0.97 0.95
udiff-l 0.95 0.93 0.66 0.83 0.82
search-replace 0.97 0.87 0.94 0.95 0.90
Gemini 2.5 Flash udiff 0.90 0.34 0.72 0.94 0.94
udiff-h 0.91 0.54 0.76 0.95 0.94
udiff-l 0.79 0.07 0.36 0.47 0.46
search-replace 0.95 0.85 0.88 0.94 0.87
@@
udiff-h 0.01 0.01 0.00 0.03 0.02
udiff-l 0.00 0.01 0.00 0.00 0.00
search-replace 0.00 0.00 0.00 0.00 0.00
Qwen2.5-Coder-1.5B udiff 0.16 0.04 0.00 0.08 0.09
udiff-h 0.16 0.05 0.01 0.09 0.11
udiff-l 0.11 0.03 0.00 0.00 0.00
search-replace 0.22 0.08 0.20 0.36 0.37
Qwen2.5-Coder-3B udiff 0.38 0.17 0.01 0.18 0.21
udiff-h 0.41 0.20 0.02 0.24 0.21
udiff-l 0.34 0.13 0.00 0.02 0.04
search-replace 0.41 0.29 0.14 0.27 0.23
Qwen2.5-Coder-7B udiff 0.58 0.65 0.06 0.42 0.34
udiff-h 0.57 0.64 0.05 0.32 0.26
udiff-l 0.57 0.62 0.00 0.00 0.00
search-replace 0.75 0.71 0.28 0.61 0.55
Qwen2.5-Coder-32B udiff 0.84 0.87 0.23 0.74 0.64
udiff-h 0.83 0.86 0.17 0.50 0.44
udiff-l 0.83 0.82 0.01 0.02 0.02
search-replace 0.92 0.85 0.68 0.91 0.82
The results in Table 5 show that search-replace is the most effective representation overall, especially for larger models, achieving strong results across all three tasks. Structured udiff and udiff-h formats remain reliable but are consistently outperformed by search-replace. In contrast, udiff-l performs poorly across most settings, suggesting that verbose line markers make the format harder for model rather than easier. Overall, search-replace offers a good balance between generation simplicity and faithful application for high-capacity models, while smaller open models still struggle regardless of representation.
@@
    Marker collisions. Small models may confuse single-character udiff markers (+, -, leading space) with ordinary code characters. Replacing them with explicit tags (ADD/DEL/CON) in udiff-l makes the control tokens unambiguous and rare; the model mainly needs to decide, for each line, which tag to use and then emit the line, an easier decision for weaker models. Nevertheless, our results suggest that this verbosity increases complexity rather than actually helping the models.
@@ 7 Related Work
A long line of benchmarks evaluate LLMs on code generation from natural language. humaneval and mbpp introduced HumanEval and MBPP, which remain the most widely used, with extensions such as HumanEval+ and MBPP+ [evalplus], HumanEval-XL [humaneval-xl], and multilingual variations such as MultiPL-E [multiple]. BigCodeBench [bigcodebench] emphasizes more complex and library-heavy tasks. These benchmarks primarily test function-level synthesis and correctness, but do not address how edits are represented or applied.
@@
More recent work has evaluated language models in editing settings. canitedit introduces instruction-following edits, and codeeditorbench evaluate debugging, polishing and translation tasks. At a larger scale, swebench proposed SWE-bench, which tasks models with resolving real GitHub issues by producing patches that apply and pass tests. These benchmarks reflect realistic workflows, but include many factors, from retrieval and long context reasoning to semantic correctness and patch formatting. These benchmarks capture realistic workflows, including patch application, but their evaluation mixes retrieval, long-context reasoning, semantic correctness, and patch formatting — making it difficult to isolate the role of edit representation.
@@
Despite this breadth, existing benchmarks all assume a fixed diff format, typically unified diff. In reality, there are multiple representations in use: normal diff, context diff, unified diff, OpenAI’s V4A diff format [gpt4.1], and search/replace schemes, including anchored replacements used by some agents. Format is therefore a real variable, not a constant. Aider’s benchmarks [aider-polyglot] are notable for reporting differences between formats (whole file rewrite, search/replace, unified diff), but their evaluation was motivated by tool design choices rather than a systematic study of edit representations.